40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu May 31 23:12:17 2018, gautam, Update, LSC, DRMI locking attempt REFL55_whtCheck.pdf
    Reply  Fri Jun 1 01:22:50 2018, gautam, Update, LSC, DRMI locking restored DRMI_20180531.png
       Reply  Fri Jun 8 01:58:56 2018, gautam, Update, LSC, DRMI locking attempt again 
          Reply  Tue Jun 12 03:22:25 2018, gautam, Update, LSC, AUX laser shuttered 
             Reply  Wed Jun 13 01:02:40 2018, gautam, Update, LSC, Reliable and repeatable 1f DRMI locking 
                Reply  Wed Jun 13 11:17:40 2018, gautam, Update, LSC, PRCL loop shape anomaly PRCL_12Jun2018_WeirdShape.pdf
                   Reply  Thu Jun 14 00:40:42 2018, gautam, Update, LSC, PRCL loop shape anomaly 
                Reply  Thu Jun 14 18:09:24 2018, gautam, Update, LSC, Reliable and repeatable 1f DRMI locking DRMI1f_June14.pdf
                   Reply  Fri Jun 15 00:53:21 2018, gautam, Update, LSC, Calibrated MICH spectrum C1NB_disp_40m_MICH_NB_2018-06-14.pdf
Message ID: 13969     Entry time: Fri Jun 15 00:53:21 2018     In reply to: 13966
Author: gautam 
Type: Update 
Category: LSC 
Subject: Calibrated MICH spectrum 

Using the numbers from the sensing measurement, I calibrated the measured in-loop MICH spectrum from Tuesday night into free-running displacement noise. For convenience, I used the noise-budgeting utilities to make this plot, but I omitted all the technical noise curves as the coupling has probably changed and I did not measure these. The overall noise seems ~x3  higher everywhere from the best I had last year, but this is hardly surprising as I haven't optimized anything for low noise recently. To summarize:

  • DRMI was locked using 1f error signals.
  • MICH was controlled using AS55_Q.
  • Main difference is that we have a little less (supposedly 10%) light on the AS55 PD now because of the AUX laser injection setup. But the AUX laser was shuttered.
  • 1f LSC PDs (REFL11, REFL55 and AS55) had ADC whitening filters engaged in while this data was taken.
  • ITM and BS coils were not de-whitened.

I will do a more thorough careful characterization and add in the technical noises in the coming days. The dominant uncertainty in the sensing matrix measurement, and hence this free-running noise spectrum, is that I haven't calibrated the actuators in a while.

Quote:

I finally analyzed the sensing measurement I ran on Tuesday evening. Sensing responses for the DRMI DOFs seems consistent with what I measured in October 2017, although the relative phasing of the DoFs in the sensing PDs has changed significantly. For what it's worth, my Finesse simulation is here

Attachment 1: C1NB_disp_40m_MICH_NB_2018-06-14.pdf  297 kB  Uploaded Fri Jun 15 01:57:27 2018  | Hide | Hide all
C1NB_disp_40m_MICH_NB_2018-06-14.pdf
ELOG V3.1.3-